mhchen:

Find 8 3x3 matricies such that A^2 = I

mhchen:

What I have so far: Let A = \[\left[\begin{matrix}a & b & c\\ d & e & f \\ g & h & i\end{matrix}\right]\] Then A^2 gives us the following system of equations: 1 = a^2 + bd + ci = e^2 + bd + fh = i^2 + cg + fh 0 = ab + be + ch = ac + bf + ci = ad + de + fg = cd + ef + fi = ag + dh + ci = bg + eh + hi We can set a,e,i to 1 and -1 (since -1^2 = 1) \[\left[\begin{matrix}1 & 0 & 0\\ 0 & 1 & 0 \\ 0 & 0 & 1\end{matrix}\right]\] \[\left[\begin{matrix}-1 & 0 & 0\\ 0 & -1 & 0 \\ 0 & 0 & -1\end{matrix}\right]\] now I'm stuck finding 6 more

mhchen:

The textbook's hint is "Look for solutions in which all entries off the main diagonal are zero"

mhchen:

Wait.....I can just do something like: (a,e,i) then all of these work: (1,1,-1) (1,-1,1) (-1,1,1) (1,-1,-1) (-1,-1,1) (-1,1,-1) And that's 6 more solutions lol.